Kanchan dyes dresses. She had to dye 30 dresses. She has so far finished 20 dresses. What fraction of dresses has she finished?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Total number of dresses Kanchan has to dye = 30 dresses
Number of dresses she has finished = 20 dresses
∴Required Fraction = 20 / 30 = 2 / 3

Thanks